1. Design a buffer that has a pH of 4.48 using one of the weak base/conjugate acid systems shown below.

Weak Base Kb Conjugate Acid Ka pKa

CH3NH2

4.2 x 10-4

CH3NH3+

2.4 x 10-11

10.62

C6H15O3N

5.9 x 10-7

C6H15O3NH+

1.7 x 10-8

7.77

C5H5N

1.5 x 10-9

C5H5NH+

6.7 x 10-6

5.17



How many grams of the chloride salt of the conjugate acid must be combined with how many grams of the weak base, to produce 1.00 L of a buffer that is 1.00 M in the weak base?

grams chloride salt of conjugate acid =

grams weak base =

2. Design a buffer that has a pH of 3.54 using one of the weak acid/conjugate base systems shown below.

Weak Acid Conjugate Base Ka pKa

HC2O4-

C2O42-

6.4 x 10-5

4.19

H2PO4-

HPO42-

6.2 x 10-8

7.21

HCO3-

CO32-

4.8 x 10-11

10.32



How many grams of the sodium salt of the weak acid must be combined with how many grams of the sodium salt of its conjugate base, to produce 1.00 L of a buffer that is 1.00 M in the weak base?

grams of sodium salt of weak acid =

grams sodium salt of conjugate base =
Answer #1

1) The base that has a pKa closest to the necessary pH is chosen, it is C5H5N.

Calculate the concentration ratio between salt and acid, using the clear equation of henderson hasselbach:

[Salt] / [Acid] = 10 ^ (pH-pKa) = 10 ^ (4.48 - 5.17) = 0.204

The acid concentration is calculated:

[Acid] = 1 M / 0.204 = 4.9 M

The grams of acid salt and base are calculated:

g Acid = M * V * MM = 4.9 M * 1 L * 115.56 g / mol = 566.24 g

g Base = 1 M * 1 L * 79.1 g / mol = 79.1 g

2) Same procedure above, choose HC2O4-, salt / acid ratio:

[Salt] / [Acid] = 10 ^ (3.54 - 4.19) = 0.224

[Salt] = 1 M * 0.224 = 0.224 M

The grams are calculated:

g Acid = 1 M * 1 L * 112.01 g / mol = 112.01 g

g Base = 0.224 M * 1 L * 111 g / mol = 24.86 g

If you like the answer, please rate it in a positive way, you will help me a lot, thank you.